OpenStudy (anonymous):

Right triangle ABC (see top of right-hand column) has one leg of length 5 in. and a hypotenuse of length 13 in. • a) Determine the length of the other leg. • b) Determine the perimeter of the triangle. • c) Determine the area of the triangle.

OpenStudy (anonymous):

a 12 b 30 c 5 * 6 = 30

OpenStudy (anonymous):

can you tell me how you got the answers?

OpenStudy (anonymous):

ok a pythagoras : x^2 = 13^2 - 5^2 = 144 x = 12 b 12+ 5 + 13 = 30 c area = 1/2 * base * height = 1/2 * 12* 5 = 30

OpenStudy (anonymous):

i din.t really do it this way because this is a 'pythagorean triad' - ie one of the right angled triangles with sides which are whole numbers the first is 5,4,3 and next is 13,12,5 but my last post is general way to do these

OpenStudy (anonymous):

confession is good for the soul!!!

OpenStudy (anonymous):

do you need any further explanation?

OpenStudy (anonymous):

no.. thank you..:)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

oh - plz click on good answer to close the question